OpenStudy (jenniferjuice):

Given the parent function of f(x) = x^3, what change will occur when the function is changed to f(x) - 3? Shift to the right 3 units Shift to the left 3 units Shift up 3 units Shift down 3 units

OpenStudy (australopithecus):

http://www.purplemath.com/modules/fcntrans.htm

OpenStudy (australopithecus):

see that link

OpenStudy (jenniferjuice):

wait so its b ?

OpenStudy (australopithecus):

https://www.youtube.com/watch?v=3rnpLI8rVb4 this video is probably better

OpenStudy (australopithecus):

f(x) - 3 = x^3 - 3

OpenStudy (australopithecus):

You should consider getting some graphing software this stuff is a lot easier when you can manipulate functions, if you are on a mac try grapher

OpenStudy (australopithecus):

Or you can just use wolframalpha

OpenStudy (jenniferjuice):

i think i can rule out a and b ?

OpenStudy (australopithecus):

There are set rules for translation and transformations just use a search engine and you should be able to find them, probably best to memorize them.

OpenStudy (australopithecus):

or you can just play with graphs by yourself and figure them out for yourself by just modifying them and looking at how they change. Just use wolframalpha if you dont have grapher, http://www.wolframalpha.com/input/?i=graph+x^3+and+graph+x^3+-+3

OpenStudy (australopithecus):

look at how it changes

OpenStudy (australopithecus):

understand why this is the case, if you set x = 0 in the function x^3 -3 you get y = -3 so you get the point (0,3) conversely, when x = 0 for x^3 you get the point, (0,0)

OpenStudy (jenniferjuice):

I DO NOT UNDERSTANDDDDDDDDD